OpenStudy (anonymous):

x + ky = 1 kx + y = 1 How do you find the value(s) of k that give this system (a) no solution, (b) a unique solution, and (c) infinitely many solutions?

OpenStudy (anonymous):

Interesting.... the identity result ... 1=1 is a infinite solutions the contadiction result... 0=1 is a no solution the exact answer... x=1 is a unique solution how do we mess with this to get each one?

OpenStudy (anonymous):

if k=1 x+y=1 x+y=1 invert one of them to -x-y=-1 and add them x+y=1 -x-y=-1 0=0 is the identity... so k=1 is an infinite solutions condition.

OpenStudy (anonymous):

what if k=-1?

OpenStudy (anonymous):

Then there's no solution?

OpenStudy (anonymous):

(for k=-1)

OpenStudy (anonymous):

correct... 0=2.. no solution

OpenStudy (anonymous):

Interesting. Hopefully I can figure out the harder problems from here. Thanks for the help!

OpenStudy (anonymous):

no problem... it looks like any other non zero number will get x=-y, which is one unique.... so I guess we are done.